A zener diode voltage regulator operated in the range 120V -180V produces a constant supply of 110V and 250mA to the load if the maximum current is equally shared between the load and zener diode then the value of load resistance and series resistance
Rl= 440 ohm , RS= 140 ohm
1
Expert's answer
2019-04-29T09:35:56-0400

Since the zener diode voltage is equal to the produced constant output supply voltage of 110 V, the load resistance can be determined according to Ohm's law:


"R_L=V_Z/I_L=110/0.250=440\space\Omega."


On the one hand, the maximum series current is


"I_{Smax}=(V_{max}-V_Z)\/R_S."

On the other, half of this maximum current equals the maximum load current according to the condition of the problem.


"\frac{1}{2}I_{Smax}=I_{Lmax},"

"I_{Smax}=0.5\text{ A}."

Hence


"R_S=(V_{max}-V_Z)/I_{Smax}=(180-110)/0.5=140\space\Omega."